OpenStudy (anonymous):

How would you evaluate the expression sin^-1(12/13)?

OpenStudy (anonymous):

with a calculator is this the actual question, or is there something else?

OpenStudy (anonymous):

for example, is the real question \[\cos\left(\sin^{-1}(\frac{12}{13})\right)\]?